# Optimal Convex Cover as Collision-free Space Approximation for Trajectory Generation

### v2

@@ Abstract
We propose an online iterative algorithm to find an optimal convex cover to under-approximate the free space for autonomous navigation to delineate Safe Flight Corridors (SFC). The convex cover consists of a set of polytopes such that the union of the polytopes represents obstacle-free space, allowing us to find trajectories for robots that lie within the convex cover. In order to find the SFC that facilitates optimal trajectory generation, we iteratively find overlapping polytopes of maximum volumes that include specified waypoints initialized by a geometric or kinematic planner. Constraints at waypoints appear in two alternating stages of a joint optimization problem, which is solved by a method inspired by the Alternating Direction Method of Multipliers (ADMM) with partially distributed variables. We validate the effectiveness of our proposed algorithm using a range of parameterized environments and show its applications for two-stage motion planning.
@@ I Introduction
All motion and path planning algorithms require the representation of the environment for collision avoidance and checking. Different algorithms choose different representations. Exact methods map obstacles from the workspace into the configuration space, this gives guarantees of completeness and optimality but may be computationally intractable [1]. Sampling-based methods do not need to represent the environment explicitly but they only offer probabilistic guarantees and may still require an excessively large number of samples in cluttered high-dimensional spaces [2]. A middle ground has become popular over the past few years, including convex approximation of the collision-free space. This is computationally attractive because it allows us to convert motion planning into an efficient optimization problem. This problem is also tractable because collision-free space can be described by a set of linear constraints corresponding to each of the polytopes in the cover.
@@
One of the key issues in this approach pertains to computing a good approximation of the environment. One could find a convex cover by a collection of convex sets such that the union of the convex sets can cover a given region or include a given guess for a trajectory. Solving the subproblem of extracting a single convex decomposition from a seed point or convex body is usually involved in generating the convex cover [3, 4, 5, 6, 7]. Considering the example in Fig. 1(a), an ideal single convex decomposition should capture most of the free space. There should also be enough overlaps between different polytopes (ellipsoids are another popular choice) in the convex cover, at least successive polytopes as a “Safe Flight Corridor” (SFC). The properties of the SFC will influence the ultimate geometric feasible region for optimization. Many existing approaches generate an SFC by selecting seed points around the initial path [8, 9, 10]. However, such covers are not ideal in practice because the selection of SFC is greedy and does not account for trajectory planning. It is difficult to determine how the geometric shapes and volumes of the polytopes interplay with the dynamical feasibility and optimality of the trajectories computed by the lower-level optimization procedure. Fig. 1(b)-(d) shows some examples of potential factors like polytope volumes and overlapping areas that potentially change the final cover of a trajectory. As a consequence, existing methods cannot give any guarantees on the quality of the SFC or the trajectory. This suggests that coupled program: front-end path planning, free space approximation using a convex cover, and trajectory generation within this cover, cannot be solved independently.
Figure 2: SFC generation pipeline in two-stage planning modules of previous works and our method. In the proposed method, we jointly optimize the path represented by waypoints together with polytopes and ellipsoids that form a convex cover of the path in free space to generate a better back-end trajectory downstream.
@@
To address optimal SFC generation, we introduce a novel algorithm that optimizes the convex cover of the environment to aid in trajectory planning for quadrotors (and other mobile robots with dynamic constraints). It updates the polytopes that form the convex cover and maximizes the volumes of ellipsoids inscribed within these polytopes given the waypoints updated by using heuristic functions to approximate trajectory cost. These include two alternating stages of a joint optimization problem, which is solved by a method inspired by the Alternating Direction Method of Multipliers (ADMM) [11] with partially distributed variables. The pipeline is shown in Fig. 2 in comparison to previous works. This approach is validated using a large variety of synthetic environments to demonstrate its flexibility and robustness.
@@
Contributions: (1) We formulate an SFC generation problem to identify the optimal convex cover in collision-free space for trajectory optimization. (2) We design a novel iterative algorithm based on the distributed ADMM for computational efficiency. (3) We extensively evaluate our method, benchmark it in environments with varying complexities, and present the first thorough analysis of the influence of the front-end plan path and SFC on the trajectory cost.
@@ II Related works
In real-world representations of geometric environments, obstacles or objects can be decomposed into a series of convex shapes to assist obstacle avoidance strategies. There are traditional methods to solve exact and approximate convex decomposition [12, 13, 14] or neural networks to learn the convex primitives [15]. However, decomposing obstacles would result in a large number of small pieces that can lead to significant computational costs. Instead, extracting the convex representation of nonconvex free space is more efficient for motion planning problem formulation [16, 17, 10, 18, 19, 20, 21]. Collision-free corridors or graphs can be generated using boxes[16, 17], sphere or ellipsoids[19, 18] and polytopes[10, 20] within environments represented by noisy point clouds or obstacle maps. A more detailed reviews of free-space decomposition are listed as follows:
@@
Single Convex Decomposition of Free Space. To extract convex regions from nonconvex environments, [3] proposed an algorithm of Iterative Regional Inflation by Semidefinite Programming (IRIS) to generate maximum volume polytopes around convex obstacles given a seed point. The iterative optimization process involved sequentially quadratic programming to find separating hyperplanes toward obstacles, followed by semidefinite programming to determine the maximum-volume ellipsoid (inner Löwner-John Ellipsoid [22]) within this polytope. However, considering the expensive computation of this approach, more works focused on improving the efficiency of the process, such as generating polytopes on line segments [8], or convex bodies [7]. With given voxel maps as inputs, [21] proposed parallel convex inflation of axis-aligned cubes on axis directions by checking the voxel points incrementally, which can achieve resolution near real-time. Moreover, in [23], the authors introduced a stereographic projection technique to project obstacle points or point clouds and convert them into a convex hull. The generative region is over-conservative for application, which is improved by [6] using sphere flipping to generate star-convex shapes and further refined into convex.
Convex Cover of a Global Map. While previous works exist for inflating a single convex region within naturally nonconvex environments, determining an optimal cover of local or entire free space remains challenging. The performance of IRIS-based decomposition for global convex covers relies heavily on the distribution of seed points, which is manually set in [3, 24]. To improve the sampling efficiency, the work [5] leveraged the topology property of the objects and employed skeleton structures to position convex primitives. Rejection sampling methods are utilized in [18, 10] to refine seed point selection outside existing polytopes or ellipsoids until the desired environment coverage is achieved. Furthermore, [20] iteratively conducted sampling to update clique covers of visibility graphs and inflate polytopes. With the convex free-space maps, the trajectory generation problem can be addressed by decoupled [10] or joint optimization [24, 25].
@@
Convex Cover along a Trajectory. For real-world deployments, computation efficiency is more important and is usually achieved by narrowing the solution space of trajectory optimization. A two-stage planning framework, involving front-end path planning with geometric constraints followed by back-end optimization within the same homotopy class, has been extensively discussed in [26] and applied to various environments decomposed by SFC. The authors in [8] proposed a jump point search (JPS) to find a geometric path composed of line segments and initialize ellipsoids on lines with one iteration polytope inflation. For general polytopes generation using seed points or lines, different paths from front-end pathfinding can be applied by iteratively checking the farthest point along the path within the polytope as the next sample points [9, 10]. However, the above methods either choose line segments or employ a greedy approach to find the next waypoint as seeds, potentially resulting in suboptimal trajectories, particularly in narrow hallways. Our approach lies within this category and focuses on decomposing collision-free space to SFC. This involves factors such as convex approximation accuracy, division of trajectory segments, and ensuring connectivity of convex regions between waypoints.
@@ III-A Trajectory Optimization Formulation
We consider the problem of finding an optimal trajectory for a $s-$order integrator system, represented by states $x(t)$ and control inputs $u(t)$, with traversal time $T$ that passes through collision-free space $\mathcal{F}_{free}$. Given a start state $x_{0}\in\mathcal{F}_{free}$ and a feasible goal region $\mathcal{F}_{goal}\subseteq\mathcal{F}_{free}$, the trajectory optimization problem is formulated as
@@ III-B Convex Cover Generation
We consider n-dimensional geometric space represented as $\mathcal{X}=\mathcal{X}_{free}\cup\mathcal{X}_{obs}\subseteq\mathbb{R}^{n}$, where $n\in\{2,3\}$. $\mathcal{X}$ is composed of free space $\mathcal{X}_{free}$ and obstacles $\mathcal{X}_{obs}$. $\mathcal{X}_{obs}$ is specified as a union of convex bodies in the set$\{\mathcal{O}_{i},i=1,2,\cdots,S\}$. The map boundary is also considered as an obstacle. Several different representations exist for the convex bodies themselves. Typically they are either convex polytopes (bounded intersections of closed half-spaces) or spheres which are commonly used while manipulating raw point-cloud observations of the environment.
@@
With such a collision-free path, for example, output by a path planning algorithm, the convex cover problem can be formulated as finding a sequence of overlapping convex sets in collision-free space that covers the given path.
In order to obtain a parameterized family of convex sets suitable for optimization, we begin by addressing the separation of obstacle space and collision-free space. As discussed in [27], according to the Supporting Hyperplane Theorem, there exists a supporting hyperplane associated with a boundary point $o_{j}$ on an obstacle $\mathcal{O}_{j}$, defined as
$\mathcal{H}_{j}=\{x\in\mathbb{R}^{n}|\ a_{j}^{T}(x-o_{j})\leq 0\}.$ (3)
We define a collision-free convex polytope as $\mathcal{P}:=\bigcap_{j=1}^{v}\mathcal{H}_{j}\subseteq\mathcal{X}_{free}$, where a subset of $v$ hyperplanes is used to delimit the polytope. A more convenient way of writing the polytope above is $\mathcal{P}=\{x\in\mathbb{R}^{n}|Ax\leq b\}$, where $A\in\mathbb{R}^{v\times n}$ and $b\in\mathbb{R}^{n}$. With regard to the previous relation, the $j$-th row of $A$ is $a_{j}^{T}$, while the $j$-th entry of $b$ is $a_{j}^{T}o_{j}$. Finding a single polytope to effectively separate obstacles from a geometric path may not always be practical, or even possible. Hence, we employ a sequence of overlapping polytopes as a SFC for covering a (preferably large) subset of the class of paths homotopic to the original path in collision-free space. The convex cover problem is to find an optimal SFC such that
@@
$\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\neq\varnothing,\ \ \forall 1\leq i\leq M-1$ (4c)
In the previous relation, we represent the $i$-th polytope as $\mathcal{P}_{i}=\{A_{i},b_{i}\}$, where $A_{i}=[a_{i,1}^{T},\cdots,a_{i,v_{i}}^{T}]^{T},b_{i}=[b_{i,1},\cdots,b_{i,v_{i}}]^{T}$. $f(\cdot)$ defines an objective function to find the best SFC. Examples of possible criteria encoded by $f$ include the volume of polytopes $f(\mathcal{P}_{i})=-vol(\mathcal{P}_{i})$, or the volume of the inner Löwner-John Ellipsoid of each polytope. Nevertheless, while previous classes of criteria are intuitively appealing, it is not obvious how they lead toward synthesizing better trajectories. In the rest of the paper, we will discuss the factors and properties of SFC that are essential for effective trajectory generation.
@@ III-C Trajectory Optimization with SFC
where $\bar{p}_{0},\bar{p}_{i},\bar{p}_{f}\in\mathbb{R}^{n\times s}$ are initial, intermediate and end derivative vectors, $\sigma_{\cdot}^{[s-1]}=(\sigma_{\cdot}(t),\cdots,\sigma_{\cdot}(t)^{(s-1)})$. Trajectory optimization with piece-wise polynomial reduces the complexity of dynamics and scaling issues of discretizations, whereas introducing nonlinearity in objectives, and constraints coupled with coefficients, times, and SFC variables. Previous research in piece-wise trajectory with corridors mainly focuses on time allocation [28, 29, 30], while the optimality and feasibility of convex approximation of environments are rarely discussed.
@@ IV Proposed Algorithms
$\min_{L_{i},d_{i},p_{i},\forall i}$ $\ w_{v}f(L_{i})+w_{c}\tilde{J}(p),\ \qquad\quad$ (7a)
@@
$h_{1}(p,\mathcal{P})\leq 0,$ (7d)
$h_{2}(L,d,\mathcal{P})\leq 0,$ (7e)
where objectives are decoupled into ellipsoids-related cost $f(L)$ and an approximation of trajectory cost $\tilde{J}(p)$ weighted by $w_{v}$ and $w_{c}$ respectively. The constraints (7b)(7c) encode the requirement that every pair of consecutive waypoints lie in the corresponding ellipsoid. To rewrite the latter pair of constraints in a more streamlined way, we introduce the function $h_{0}(\mathcal{E},p)=||L^{-T}(p-d)||_{2}-1$. The functions $h_{1}(\cdot),h_{2}(\cdot)$ represent independent constraints for ellipsoids and trajectory, with polytopes denoted as $\mathcal{P}=(\mathcal{P}_{1},\cdots,\mathcal{P}_{M})$. We use the notation $\mathcal{E}=(\mathcal{E}_{1},\cdots,\mathcal{E}_{M})$, where $\mathcal{E}_{i}=\{L_{i},d_{i}\}$, and rewrite the problem as
$\vspace{-0.3cm}\min_{\mathcal{E},p}$ $\ w_{v}\sum_{i=1}^{M}f(\mathcal{E}_{i})+w_{c}\tilde{J}(p),$ (8a)
@@
$h_{1}(p,\mathcal{P})\leq 0,$ (8d)
$h_{2}(\mathcal{E}_{i},\mathcal{P})\leq 0,\ \forall i=1,\cdots,M,$ (8e)
where $g(\cdot)$ is a relaxed equality function, for example, $g(\cdot)=\max\{0,\cdot\}^{2}$ for coupled conditions of ellipsoid $\mathcal{E}$ and waypoint $p$ variables. The independent constraints (8d) and (8e) are complex and may introduce additional variables such as trajectory time intervals and polytopes. Therefore, we incorporate the coupled constraints and formulate the scaled-form augmented Lagrangian as
$\mathcal{L}(\mathcal{E},p,y)$ $=w_{v}\sum_{i=1}^{M}f(\mathcal{E}_{i})+w_{c}\tilde{J}(p)$ (9)
$+\frac{\rho}{2}\sum_{i=1}^{M}\|[g(h_{0}(\mathcal{E}_{i},p_{i})),g(h_{0}(\mathcal{E}_{i},p_{i-1}))]^{T}$
$+y_{i}\|_{2}^{2}-\frac{\rho}{2}\|y\|_{2}^{2},$
where $y=(y_{1},\cdots,y_{M})$ are Lagrange multipliers, $\rho>0$ is a parameter.
We propose an iterative algorithm to address the problem, described as Alg. (1). A method based on a scaled-form distributed ADMM approach is applied with the relaxation of the independent constraints. The inputs include an initial geometric path and obstacles, while the output is SFC represented by polytopes. The constraints $h(\cdot)$ would be satisfied during each sub-problem. The ellipsoids’ cost function and constraints are separable and can be updated in parallel.
@@
             $p^{k+1}:=\operatorname*{arg\,min}_{p}\{\mathcal{L}(\mathcal{E}^{k},p,y^{k})|\ h_{1}(p,\mathcal{P})\leq 0\}$
@@
Algorithm 1 Given an initial geometric path and obstacles, find optimal SFC to cover the path while avoiding all the obstacles.

### v3

# Towards Optimizing a Convex Cover of Collision-Free Space for Trajectory Generation

@@ Abstract
We propose an online iterative algorithm to optimize a convex cover to under-approximate the free space for autonomous navigation to delineate Safe Flight Corridors (SFC). The convex cover consists of a set of polytopes such that the union of the polytopes represents obstacle-free space, allowing us to find trajectories for robots that lie within the convex cover. In order to find the SFC that facilitates trajectory optimization, we iteratively find overlapping polytopes of maximum volumes that include specified waypoints initialized by a geometric or kinematic planner. Constraints at waypoints appear in two alternating stages of a joint optimization problem, which is solved by a novel heuristic-based iterative algorithm with partially distributed variables. We validate the effectiveness of our proposed algorithm using a range of parameterized environments and show its applications for two-stage motion planning.
@@ I Introduction
All motion and path planning algorithms require the representation of the environment for collision avoidance. Different algorithms choose different representations. Exact methods map obstacles from the workspace into the configuration space, this gives guarantees of completeness and optimality but may be computationally intractable [1]. Sampling-based methods do not need to represent the environment explicitly but they only offer probabilistic guarantees and may still require an excessively large number of samples in cluttered high-dimensional spaces [2]. A middle ground has become popular over the past few years, including convex approximation of the collision-free space. This is computationally attractive because it allows us to convert motion planning into an efficient optimization problem. This problem is also tractable because collision-free space can be described by a set of linear constraints corresponding to each of the polytopes in the cover.
@@
One of the key issues in this approach pertains to computing a good approximation of the environment. One could find a convex cover [3] by a collection of convex sets such that the union of the convex sets can cover a subset of the collision-free region or include a given guess for trajectories. Solving the subproblem of extracting a single convex region from a seed point or convex body is usually involved in generating the convex cover [4, 5, 6, 7, 8]. Considering the example in Fig. 1(a), an ideal single convex region should capture most of the free space. There should also be enough overlaps between different polytopes (ellipsoids are another popular choice) in the convex cover, at least successive polytopes as a “Safe Flight Corridor” (SFC). The properties of the SFC will influence the ultimate geometric feasible region for optimization. Many existing approaches generate an SFC by selecting seed points around the initial path [9, 10, 11]. However, such covers are not ideal in practice because the selection of SFC is greedy and does not account for trajectory planning. It is difficult to determine how the geometric shapes and volumes of the polytopes interplay with the dynamical feasibility of the trajectories computed by the lower-level optimization procedure. Fig. 1(b)-(d) shows some examples of potential factors like polytope volumes and overlapping areas that can potentially change the final trajectory. As a result, the existing pipelines (illustrated at the top part of Fig. 2) cannot give any guarantees on the quality of the SFC or the trajectory. This suggests that coupled program: front-end path planning, free space approximation using a convex cover, and trajectory generation within this cover, cannot be solved independently.
Figure 2: SFC generation pipeline of previous works and our method. Given waypoints from the path planning, the SFC generator selects seed points or lines from waypoints and outputs overlapping polytopes. The proposed method novelly formulates this process as a joint optimization problem to generate a better back-end trajectory downstream.
@@
To address the problem of SFC generation, we introduce a novel algorithm that optimizes the convex cover of the environment to aid in trajectory planning for quadrotors (and other mobile robots with dynamic constraints). It updates the polytopes that form the convex cover and maximizes the volumes of ellipsoids inscribed within these polytopes given the waypoints updated by using heuristic functions to approximate trajectory cost. These include two alternating stages of a joint optimization problem, which is solved by a heuristic iterative approach with partially distributed variables, inspired by the Alternating Direction Method of Multipliers (ADMM) [12] with partially distributed variables. The pipeline is shown in Fig. 2 in comparison to previous works. This approach is validated using a large variety of synthetic environments to demonstrate its flexibility and robustness.
@@
Contributions: (1) We formulate an SFC generation problem to optimize a convex cover of a subset of the collision-free space. This subset is approximated by the boundaries of trajectories that belong to the same homotopy class as a given path within the collision-free space for trajectory optimization. (2) We design a novel iterative algorithm based on heuristics with partially distributed variables to maximize the efficiency of the computation. (3) We extensively evaluate our method, benchmark it in environments with varying complexities, and present the first thorough analysis of the influence of the front-end plan path and SFC on the trajectory cost.
@@ II Related works
In real-world representations of geometric environments, obstacles or objects can be decomposed into a series of convex shapes to assist obstacle avoidance strategies. There are traditional methods to solve exact and approximate convex region [13, 14, 15] or neural networks to learn the convex primitives [16]. However, decomposing obstacles would result in a large number of small pieces that can lead to significant computational costs. Instead, extracting the convex representation of nonconvex free space is more efficient for motion planning problem formulation [17, 18, 11, 19, 20, 21, 22]. Collision-free corridors or graphs can be generated using boxes[17, 18], sphere or ellipsoids[20, 19] and polytopes[11, 21] within environments represented by point clouds or obstacle maps.
@@
To extract convex regions from nonconvex environments, [4] proposed an algorithm of Iterative Regional Inflation by Semidefinite Programming (IRIS) to generate maximum volume polytopes around convex obstacles given a seed point. The iterative optimization process involved sequentially quadratic programming to find separating hyperplanes toward obstacles, followed by semidefinite programming to determine the maximum-volume ellipsoid (inner Löwner-John Ellipsoid [23]) within this polytope. However, considering the expensive computation of this approach, most other works focused on improving the efficiency of the process, such as generating polytopes on line segments [9], or convex bodies [8]. With given voxel maps as inputs, [22] proposed parallel convex inflation of axis-aligned cubes on axis directions by checking the voxel points incrementally, which can achieve resolution near real-time. Moreover, in [24], the authors introduced a stereographic projection technique to project obstacle points or point clouds and convert them into a convex hull. The generative region is over-conservative for application, which is improved in [7] using sphere flipping to generate star-convex shapes and further refined into convex. The work in [8] improves the polytope inflation and maximum volume inscribed ellipsoid processes in IRIS, providing a more efficient solution through a linear-complexity algorithm. Except for inflating a single convex region within naturally nonconvex environments, determining a convex cover of local or global space remains challenging. Several offline methods face challenges in selecting appropriate seed points to effectively generate a convex cover of the environment [4, 25]. To improve the sampling efficiency, the work [6] leveraged the topology property of the objects and employed skeleton structures to position convex primitives. Furthermore, [21] iteratively conducted sampling to update clique covers of visibility graphs and inflate polytopes. With the convex free-space maps, the trajectory generation can be addressed by mix integer programming [25, 26].
@@
For real-world applications to trajectory optimization, computational efficiency is crucial and is typically achieved by decoupling the problem into subproblems or by narrowing the solution space. A two-stage planning framework, involving front-end path planning with geometric constraints followed by back-end trajectory optimization within the same homotopy class, has been extensively discussed in [27] and applied to SFC-decomposed environments. The authors in [9] proposed a jump point search (JPS) to find a geometric path composed of line segments and initialize ellipsoids on lines with one iteration polytope inflation. For general polytopes generation using seed points or lines, different paths from front-end pathfinding can be applied by iteratively checking the farthest point along the path within the polytope as the next sample points [10, 11]. However, the above methods either fix line segments or employ a greedy approach to find the next waypoint as seeds, potentially resulting in suboptimal trajectories, particularly in narrow hallways.
Unlike previous methods, we introduce a novel approach that optimizes the selection of seeds in polytope generation for better trajectory optimization. Our approach uniquely focuses on decomposing collision-free space to SFC, addressing critical factors such as convex approximation accuracy, trajectory segment division, and connectivity of convex regions between waypoints. This integrated approach ensures that the trajectory optimization process is both optimized and computationally efficient, setting our method apart from existing techniques.
@@ III-A Trajectory Optimization Formulation
We consider the problem of finding an optimal energy minimization trajectory for a $s-$order integrator system, represented by states $x(t)$ and control inputs $u(t)$, with traversal time $T$ that passes through collision-free space $\mathcal{F}_{free}$. Given a start state $x_{0}\in\mathcal{F}_{free}$ and a feasible goal region $\mathcal{F}_{goal}\subseteq\mathcal{F}_{free}$, the trajectory optimization problem is formulated as
@@ III-B Convex Cover Generation
We consider n-dimensional geometric space represented as $\mathcal{X}=\mathcal{X}_{free}\cup\mathcal{X}_{obs}\subseteq\mathbb{R}^{n}$, where $n\in\{2,3\}$. $\mathcal{X}$ consists of free space $\mathcal{X}_{free}$ and obstacles $\mathcal{X}_{obs}$, where $\mathcal{X}_{obs}$ is defined as a union of convex bodies $\mathcal{O}=\{\mathcal{O}_{i}\}_{i=1}^{S}$. Several different representations exist for the convex bodies themselves. Typically they are either convex polytopes (bounded intersections of closed half-spaces) or spheres which are commonly used while manipulating raw point-cloud observations of the environment.
@@
With such a collision-free path, for example, output by a path planning algorithm, the convex cover problem can be formulated as finding a sequence of overlapping convex sets in collision-free space that covers the given path. To obtain a parameterized family of convex sets suitable for optimization, we begin by defining a collision-free convex polytope using $\mathcal{H}$-representation $\mathcal{P}(A,b)=\{x\in\mathbb{R}^{n}|Ax\leq b\}\subseteq\mathcal{X}_{free}$, with $A\in\mathbb{R}^{v\times n}$ and $b\in\mathbb{R}^{n}$. We employ a sequence of overlapping polytopes for covering a subset of the collision-free space that contains the class of paths homotopic to the original path. Hence, the convex cover of a given path is to generate an SFC such that
@@
$\mathcal{P}_{i}\cap\mathcal{P}_{i+1}\neq\varnothing,\ \ \forall 1\leq i\leq M-1.$ (3c)
We represent sequential polytopes $\mathcal{P}=(\mathcal{P}_{1},\cdots,\mathcal{P}_{M})$, with the $i$-th polytope $\mathcal{P}_{i}$ parameterized by $A_{i}=[a_{i,1}^{T},\cdots,a_{i,v_{i}}^{T}]^{T}$ and $b_{i}=[b_{i,1},\cdots,b_{i,v_{i}}]^{T}$. The function $f(\cdot)$ defines an objective function to find the best SFC. Examples of possible criteria encoded by $f$ include the volume of polytopes $f(\mathcal{P}_{i})=-vol(\mathcal{P}_{i})$, or the volume of the inner Löwner-John Ellipsoid of each polytope. Nevertheless, while previous classes of criteria are intuitively appealing, it is not obvious how they lead toward synthesizing better trajectories. In the rest of the paper, we will discuss the factors and properties of SFC that are essential for effective trajectory generation.
@@ III-C Trajectory Optimization with SFC
where $\bar{p}_{0},\bar{p}_{i},\bar{p}_{f}\in\mathbb{R}^{n\times s}$ are initial, intermediate and end derivative vectors, $\sigma_{\cdot}^{[s-1]}=(\sigma_{\cdot}(t),\cdots,\sigma_{\cdot}(t)^{(s-1)})$. As discussed in [29], the piecewise polynomial trajectory can be parameterized as $\sigma(t)=\sigma(p,\tau)(t)$, where $p=(\bar{p}_{0},\cdots,\bar{p}_{i},\cdots,\bar{p}_{f})$. For brevity, we will use $p$ and $\tau$ to represent the parameters that determine the polynomial trajectory. Trajectory optimization with piece-wise polynomial reduces the complexity of dynamics and scaling issues of discretizations, whereas introducing nonlinearity in objectives, and constraints coupled with coefficients, times, and SFC variables. Previous research in piece-wise trajectory with corridors mainly focuses on time allocation [28, 30, 31], while the optimality and feasibility of convex approximation of environments are rarely discussed.
@@ IV Proposed Algorithms
$\min_{L_{i},d_{i},p_{i},\forall i}$ $\ w_{v}\tilde{f}(L_{i})+w_{c}\tilde{J}(p,\tau),\ \qquad\quad$ (6a)
@@
$h_{1}(p,\tau,\mathcal{P})\leq 0,$ (6d)
$h_{2}(\mathcal{E}(L,d),\mathcal{P})\leq 0.$ (6e)
The objectives are decoupled into two parts: the approximation of the polytope volume in Eq. (3a) as the volume of inner ellipsoid $\tilde{f}(L)$, and an approximation of trajectory cost in Eq. (5a) as $\tilde{J}(p,\tau)$. These terms are weighted by $w_{v}$ and $w_{c}$ respectively. The constraints (6b) and (6c) encode the requirement that every pair of consecutive waypoints lie in the corresponding ellipsoid. To rewrite the latter pair of constraints in a more streamlined way, we introduce the function $h_{0}(\mathcal{E},p)=||L^{-T}(p-d)||_{2}-1$. The functions $h_{1}(\cdot),h_{2}(\cdot)$ in (6b) and (6c) represent separate conditions for ellipsoids and trajectory, which are further elaborated in section IV-B and section IV-C. We adopt the notation $\mathcal{E}=(\mathcal{E}_{1},\dots,\mathcal{E}_{M})$, where each $\mathcal{E}_{i}$ is parameterized by $L_{i}$ and $d_{i}$. We reformulate the problem as follows:
$\min_{\mathcal{E},p}$ $\ w_{v}\sum_{i=1}^{M}\tilde{f}(\mathcal{E}_{i})+w_{c}\tilde{J}(p,\tau),$ (7a)
@@
$h_{1}(p,\tau,\mathcal{P})\leq 0,$ (7d)
$h_{2}(\mathcal{E},\mathcal{P})\leq 0,$ (7e)
where $g(\cdot)=\max\{0,\cdot\}^{2}$ for coupled conditions of ellipsoid $\mathcal{E}$ and waypoint $p$ variables. The independent constraints (7d) and (7e) are complex and may introduce additional variables such as trajectory time intervals and polytopes. Therefore, we incorporate the coupled constraints and formulate the scaled-form augmented Lagrangian as
$\mathcal{L}(\mathcal{E},p,y)$ $=w_{v}\sum_{i=1}^{M}\tilde{f}(\mathcal{E}_{i})+w_{c}\tilde{J}(p,\tau)-\frac{\rho}{2}\|y\|_{2}^{2}$ (8)
$+\sum_{i=1}^{M}\frac{\rho}{2}\|[g(h_{0}(\mathcal{E}_{i},p_{i})),g(h_{0}(\mathcal{E}_{i},p_{i-1}))]^{T}+y_{i}\|_{2}^{2}$
where $y=(y_{1},\cdots,y_{M})$ are Lagrange multipliers, $\rho>0$ is a parameter. As our problem consists of decoupled subproblems, including convex region inflation, ellipsoid generation, and trajectory generation, multi-level decomposition [32, 33, 34, 12] provides a suitable approach for efficiently solving the original problem. To solve the problem, we propose an iterative algorithm, detailed in Alg. (1), that leverages a scaled-form ADMM [12] in a partially distributed manner. We relax the constraints on independent variables of subproblems for efficient updates. The input variables include an initial geometric path and obstacles, while the output is SFC represented by a sequence of polytopes. The constraints $h(\cdot)$ would be satisfied during each sub-problem. Additionally, the cost function and constraints for ellipsoid generation are separable and can be updated in parallel.
@@
             $p^{k+1}:=\operatorname*{arg\,min}_{p}\{\mathcal{L}(\mathcal{E}^{k},p,y^{k})|\ h_{1}(p,\tau,\mathcal{P})\leq 0\}$
@@
Algorithm 1 Given an initial geometric path and obstacles, find an SFC to cover the path while avoiding all the obstacles.